OpenStudy (anonymous):

Which of the following is a correct equation for the line passing through the point (-1,4) and having slope m = -3?

OpenStudy (anonymous):

Y= -3x +11

OpenStudy (anonymous):

Check all that apply answer choices A y=-3x+1. B 3x+Y=1. C y=-1/4x-3. D y-4=3(x+1)

OpenStudy (academicgurusinc):

First of all, remember what the equation of a line is: y = mx+b So (-1,4) y=mx+b or 3=-3x(-1+b) Solve for B and you'll get b=3-(-3)(-1) and B=1 So you'll get y=-3x+1 as your final answer. @Austin66 is wrong. Hope that helps. If you found this helpful, I encourage to subscribe to our youtube channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)

OpenStudy (academicgurusinc):

AUSTIN'S ANSWER IS WRONG.

OpenStudy (anonymous):

Ok thanks academicgurusinc

OpenStudy (academicgurusinc):

You're welcome :)

OpenStudy (triciaal):